OpenStudy (anonymous):

What is the answer of: 0.05(6x)+0.25(1+6x)+1(x)=11.45?

OpenStudy (anonymous):

expand it out so you can combine like terms

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

just multiply the terms. like 0.05(6x) = 0.30x

OpenStudy (anonymous):

Can youshow it to me how? i Don't get it. -_-

OpenStudy (whpalmer4):

\[0.05(6x)+0.25(1+6x)+1(x) = 11.45\] \[0.3x + 0.25 + 1.5x + 1x = 11.45\] \[2.8x + 0.25 = 11.45\] \[2.8x = 11.2\] \[x = 11.2/2.8 = 4\]

OpenStudy (anonymous):

Thanks. :)

OpenStudy (whpalmer4):

Do you see how that worked?

OpenStudy (anonymous):

Yeah, but not that much. But I'll try myself to understand it..

OpenStudy (whpalmer4):

I'm just using the property that a (b + c) = ab + ac

OpenStudy (anonymous):

ok! remember to close the question guys :)